Status Report

ISS On-Orbit Status 9 Feb 2002

By SpaceRef Editor
February 9, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below.

First day of a well-earned weekend for the crew. However, their
schedule as usual did have a few routine tasks.

After the morning DPC (daily planning conference) with Flight
Control, the crew performed the regular weekly station clean-up, a
2-hr. scrub-down of their living and working area, including
ventilation grilles.

CDR Onufrienko then performed the daily SOSH life support systems
maintenance, while FE-1 Carl Walz checked automated payloads

FE-2 Dan Bursch prepared the regular “delta” file for the IMS
(inventory management systems) for downlinking, to be incorporated as
an update in the inventory database maintained at MCC-M.

The crew had time reserved on today’s schedule for reviewing and
studying upcoming changes in the U.S. segment C&W (caution and
warning) system and PCS (portable computer system) due to the major
8A software transition next week.  More details were uplinked to
familiarize them with the transition events taking place over the
next month or so.  The procedure, which will affect the C&C
(command and control) and GNC MDMs (guidance, navigation and control
multiplexer/demultiplexers [computers]) as well as the MSS (mobile
service system, also known as SSRMS/Canadarm2), is preferably called
8A software transition instead of R2 (Release 2) transition because
the payload software (PEP) is already R2 and will be upgraded to R3.
The 1000 files required for the 8A transition are already on board9FeD
having come up in SSMMU5 (solid-state mass memory unit #5).  A
readiness review with crew participating is scheduled for 2/12, and
the 8A “mega” procedure will then begin on 2/14. It is expected to
have two C&Cs and one GNC MDM running the new s/w by the end of
lunch that day.  To simplify the manual pointing of the S-band
antenna while the GNC MDM is off-line (about 4-6 hours), the station
will maneuver to XPOP attitude for that duration.  The “mega”
procedure has been defined in 21 individual steps, ending with the
loading of the PEP R3 s/w.

Planning for the EVA planned for 2/20 is progressing, and more
details were uplinked to the crew. These include a breakdown of the
new DOUG (dynamic orbital ubiquitous graphics) package, which
provides computerized 3D representation of ten major translation path
segments to be taken by Bursch and Walz during the six-hour
excursion.  Between now and 2/20, the crew will also be able to
practice various “flying scenarios” on the hand-controller-operated
onboard trainer for the SAFER (simplified aid for EVA rescue)
emergency propulsive devices.  The PC-game-like computer model
is programmed both for executing specific scenarios with expected
translation paths and for random scenarios and random separation

The 360-deg panoramic photographs shot by the crew last week inside
the ISS as part of the new SIS (Surround Imagery for Station)
experiment have been “stitched” together on the ground, causing great
excitement for its potential for PAO purposes, training, inventory
interface, etc.  Some follow-up questions were uplinked, to help
with future planning for more SIS.  It may involve
surround-imaging of the entire station, and specialists need to
figure out how best to get the full sphere (nadir and zenith in
addition to horizontal panorama) and how many locations are needed to
get optimum coverage of each module.

After the protracted troubleshooting of the S-band LDR (low data
rate) problem of 12/26/01, the ground today scheduled a test of the
LDR channel.  For about 20 minutes during a comm pass, S-band
was moded to LDR and then back to HDR (high data rate), followed by a
voice check with the crew on HDR.

The Russian Uragan (“hurricane”) and associated Japanese HDTV (high
definition TV) experiment is usually not part of the daily work plan
but is regularly being performed by the CDR in his spare time. 
Main purpose of the current phase of Uragan is to test the new
Rubinar binocular telescope with its electronics and the quality of
LIV video sent to MCC-M.  The experiment uses a Kodak 460 DCS
(digital camera system), a Nikon F5 camera and the LIV video
system.  Today’s geographical target areas were North Thailand,
North Vietnam and China till the Yangtze River, part of Madagascar,
West coast of India, parts of the Himalayas with Mt. Everest, Tibet,
the Kilimanjaro volcano, the Bab-al-Mandeb Straits and Oman Gulf
(both with “any oil terminal”) and Kabul. For HDTV, preferred imagery
areas were India, Himalayas, Tibet, Mt. Everest, Gobi Desert and

Science Update (Expedition Four —  9th):

As the Lead Increment Scientist was able to report to the crew,
the only science impact from the loss of power during the Russian TVM
computer failure on 2/4 may be the samples in the BTR (biotechnology
refrigerator), but that is not known for sure until the Principal
Investigators (PIs) get the samples back.  The temperature did
not rise as much as was thought at first, so the ground team is
optimistic.  Thanks were extended to the crew for the quick
recovery and all their continuing excellent work on the payloads and
the Lab window survey.

Hoffman-Reflex:  In progress.

Extra-Vehicular Activity Radiation Monitors (EVARM):  The
third EVARM badge reading was successfully completed.  The next
Pre-EVA read is scheduled for 2/13.

Ultrasound:  FE-2 Bursch conducted a successful checkout
of the HRF (human research facility) Ultrasound hardware, and
researchers are happy that “he had some fun doing it”.  The
image files were received on the ground, and the team is looking
forward to evaluating the quality of the on-orbit ultrasonic

Pulmonary Function in Flight (PuFF): The PuFF team enjoyed
watching the real-time data from Saturday’s session.  Thanks to
the crew’s excellent work and some great Ku-band coverage, they saw
all of Dan’s tests and about half of Carl’s.  The team noticed
that the crewmembers are taking the time to examine their test
performance and repeat tests when necessary.  The team really
appreciates these efforts to obtain PuFF data, especially so close to
the EVA.

Renal (Kidney) Stone Experiment:  The Renal team thanked
the crew for their participation.  The team looks forward to
continuing the study.
Interactions (NTXN):  Continuing on a weekly schedule,
nominally on Wednesdays.

Human Research Facility Workstation (HRF WS):  The HRF Workstation Functional was successfully
completed. The crew was thanked for thedbefforts and patience with
the difficult cabling operation.

Human Research Facility/PC (HRF/PC):  This week’s
downlink sent 1/15 PuFF files, last week’s EVARM files, and Friday’s
Ultrasound images down to the PIs on the ground.  The PIs are
looking forward to analyzing this data..

Cellular Biotechnology Support Systems (CBOSS): The cold
storage phase of the CBOSS experiment is in progress.

Physics of Colloids in Space (EXPPCS):  Examination of
the continued growth of the fractal network in EXPPCS’ slow fractal
sample continued this week during 120 hours of operations. Most of
the measurements are dynamic light scattering measurements, which are
sensitive to the amount of motion of the colloid particles, which can
in turn be related to the size of the clusters of particles. The
results show that the clusters are growing exceedingly slowly and
probably have not yet formed a gel. [A fractal is something that
appears to have the same structure under different degrees of
magnification. One example is the coastline of a continent: maps
showing 25 miles or 250 miles of coastline will appear quite similar
in the apparent amount of roughness. The fractal material under
study, a mix of a colloid with a salt to initiate fractal gel growth
— of interest to science for its structure and behavior in a variety
of Earth-based manufacturing fields — is 99.992 percent water and
only 0.008 percent solid. Among the questions asked by scientists is
whether something with so little material will be able to form a gel,
and how does the gel age].

Space Acceleration Measurement System (SAMS):
Currently all five
sensors are actively transmitting acceleration data.  Will
continue to support general characterization in the vibratory regime
of the microgravݯ”environment.

Microgravity Acceleration Measurement System (MAMS): 
MAMS  continues collection and archival of low-frequency
acceleration data for the characterization of the ISS quasi-steady
microgravity environment. HiRAP has been  disabled due to
continuing coverage by SAMS and to reduce the downlink bandwidth
utilized by acceleration measurements.

Protein Crystal Growth-Single Locker Thermal Enclosure System
Researchers did not experience any problems with the
temperatures in the STES units as a result of the power outage on
2/4. Apparently the temperature stayed cool in the laboratory, so
that the STES units did not heat up. This is good news. 
Activation of the #7 cylinder in the 007 unit was nominal.

Materials ISS Experiment (MISSE): Nominal and collecting

Educational Payload Operations (EPO):  EPO will most like
be rescheduled for Week 11 (was cancelled on 2/4 because of the RS
computer problem and power outage).

Active Rack Isolation System – Isolation Characterization
Experiment (ARIS-ICE)
:  ARIS-ICE is completing a detailed
plan to recover from the ARIS pushrod failure and prevent future
failures.  That plan, which includes several crew activities,
will begin implementation next week.  ARIS and ARIS-ICE have
enjoyed great success to date, and the team is looking forward to
continued success once operations have started again.
EarthKAM:  All systems and software are operational and
running smoothly.  The ground team  continues to process
photo requests and post images to the Datasystem. Pictures received
so far using the 180mm lens are clear and “wonderfully detailed”. The
students are dazzled at the increased level of detail resulting from
the lens swap to the 180mm lens.  All schools have successfully
received images back, and continue to submit photo requests at an
“incredible ever-increasing rate!” Some stats: Number of images
requested (as of 2/8): 875; number of images taken and downloaded:
606; number of schools involved: 18 active, 8 offline.  All
EarthKAM images are available for public access on the EarthKAM data
system at:

Advanced Astroculture (ADVASC):  The ADVASC Cable
Connections were completed on Saturday (2/2).  ADVASC is
scheduled to be activated next Wednesday morning (2/13).

Zeolite Crystal Growth (ZCG):  ZCG team was very excited
to find out that the furnace worked as expected during the check-out
activity.  Furnace heat-up, telemetry reception, and ground
commanding were all nominal.  Hardware is ready to process
samples to be launched on ISS 8A.
Crew Earth Observations (CEO):  Optional targets today
were Somalia Coast (pass was well inland from the coast, so
crew was asked to look the right of track and map the vegetation
contrast near the coast during this dry season)
, E.
Mediterranean Dust and Smog
(a slow-moving winter storm
continues to churn in the central Mediterranean. Of interest, as ISS
crossed the Libyan coast from the SW: oblique views of possible dust
clouds being swept northward over the sea)
, Lake Poopo
(an excellent pass over this target for careful mapping of this
Altiplano basin, with near-vertical views, from the bright salars
through Lake Poopo itself, noting color changes and the presence of
, Tuamotu-Austral Islands (good light and weather
for nadir views with the ESC of reefs and atolls on the extreme
eastern end of this long, narrow archipelago. Second pass continued
northeastward over the broader, more extensive Tuamotus, requiring
use of the ESC for vertical views to map reefs and atolls there)
and Eastern United States (both illumination and weather
conditions favored views of aerosol plumes forming over the Ohio
River Valley. Crew was to try for oblique and limb shots to the right
of track).

ISS Orbit (as of this morning, 5:59 am EST):
Mean altitude — 385.8 km

  • Apogee — 389.8 km
  • Perigee — 381.7 km
  • Period — 92.3 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0005966
  • Orbits per 24-hr. day — 15.61
  • Decay rate — 370 m (mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. â98) — 18410
  • Current Flight Attitude — LVLH +XVV ZLV (local vertical/local
  • horizontal: +X-axis in velocity vector; Z-axis in local vertical),
  • with TEA (torque equilibrium attitude [pitch: -10 deg, yaw: -7 deg,
  • roll: 0 deg]).

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.